Portions of the financial statements for Peach Computer are provided below.

PEACH COMPUTER
Income Statement
For the year ended December 31, 2018
  Net sales $1,625,000
  Expenses:
      Cost of goods sold $980,000
      Operating expenses 490,000
      Depreciation expense 43,000
      Income tax expense 33,000
        Total expenses 1,546,000
  Net income $ 79,000


PEACH COMPUTER
Selected Balance Sheet Data
December 31
2018 2017 Increase (I)
or
Decrease (D)
  Cash $95,000    $81,500 $13,500 (I)
  Accounts receivable 46,700    52,500 5,800 (D)
  Inventory 68,000    51,500 16,500 (I)
  Prepaid rent 2,300    3,600 1,300 (D)
  Accounts payable 38,000    33,500 4,500 (I)
  Income tax payable 4,300    6,500 2,200 (D)


Required:

Prepare the operating activities section of the statement of cash flows for Peach Computer using the direct method. (List cash outflows as negative amounts.)
Answer #1

Statement of cash flow – Direct method

Cash flow from operating activities

Cash receipts from customers

1,630,800

Less: Cash payment to suppliers

- 992,000

Cash payment for operating expenses

- 488,700

Cash payment for income tax

- 35,200

Net cash provided by Operating Activities

114,900

Cash receipts from customers = Revenue from sales + Decrease in accounts receivables

= 1,625,000 + 5,800

= $1,630,800

Cash payment to suppliers = Cost of goods sold + Increase in inventory - Increase in accounts payable

= 980,000 + 16,500 - 4,500

= $992,000

Cash payment for operating expenses = Operating expenses - Decrease in prepaid rent

= 490,000 - 1,300

= $488,700

Cash payment for income tax = Income tax expense + Decrease in income tax payable

= 33,000 + 2,200

= $35,200

Kindly comment if you need further assistance. Thanks